2. Project S costs $10,000 and its expected cash flows would be $5,500 per year for 5 years. Mutually exclusive Project L costs $45,000 and its expected cash flows would be $10,250 per year for 5 years. If both projects have a WACC of 15%, which project would you recommend?

Select the correct answer.

a. Both Projects S and L, since both projects have NPV's > 0.
b. Both Projects S and L, since both projects have IRR's > 0.
c. Project L, since the NPVL > NPVS.
d. Neither Project S nor L, since each project's NPV < 0.
e. Project S, since the NPVS > NPVL.

3. A project has annual cash flows of $3,500 for the next 10 years and then $7,500 each year for the following 10 years. The IRR of this 20-year project is 13.67%. If the firm's WACC is 11%, what is the project's NPV? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

1.
Recommend Project S, since the NPVS > NPVL.  

NPV of S=-10000+5500/15%*(1-1/1.15^5)=8436.85303906

NPV of L=-45000+10250/15%*(1-1/1.15^5)=-10640.41024538

2.
=-(+3500/13.67%*(1-1/1.1367^10)+7500/1.1367^11*(1-1/1.1367^10)/(1-1/1.1367))+3500/11%*(1-1/1.11^10)+7500/1.11^11*(1-1/1.11^10)/(1-1/1.11)
=6669.68631020